OpenStudy (bieber896):

Check my answer?? -2 <= d + 1/2 < 4 and a 1/2

OpenStudy (bieber896):

@haleyelizabeth2017

OpenStudy (bieber896):

MY ANSWER: [2.5, 4)

OpenStudy (haleyelizabeth2017):

I'm not sure if I remember this stuff lol sorry!

OpenStudy (bieber896):

aww haha it's okay

OpenStudy (haleyelizabeth2017):

:( @Beleaguer might know :D

OpenStudy (anonymous):

Whats the question its asking you

OpenStudy (anonymous):

@Bieber896

OpenStudy (bieber896):

To solve the compound inequality

OpenStudy (haleyelizabeth2017):

\[-2 \le d+\frac{ 1 }{ 2 } < 4 \frac{ 1 }{ 2 }\] is this what it looks like?

OpenStudy (bieber896):

haha yes

OpenStudy (haleyelizabeth2017):

you got it correct but it is actually written as \(-2\frac{ 1 }{ 2 } \le d < 4\)

OpenStudy (anonymous):

Yea, I think that is right because you have to remove the 1/2 from the middle. I barely remember this, but it looks right. Good job Haley

OpenStudy (haleyelizabeth2017):

:) why thank you lol it's only been a year haha

OpenStudy (bieber896):

Haha thanks :) @Beleaguer @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

yep yep :D

OpenStudy (anonymous):

-2 <= d + 1/2 <= 4 1/2 -2 <= d + 1/2 <= 9/2 -- subtract 1/2 -2 - 1/2 <= d <= 9/2 - 1/2 -4/2 - 1/2 <= d <= 8/2 -5/2 <= d <= 4 or - 2.5 <= d <= 4 yep..you got it

OpenStudy (anonymous):

oops...there is not supposed to be an equal sign...it should be : -2.5 <= d < 4

OpenStudy (bieber896):

@kelliegirl33 but what if there are 2 variables?

OpenStudy (anonymous):

give me an example

OpenStudy (anonymous):

oh crap..I lost track of time. Sorry....I will be back on later

OpenStudy (bieber896):

5m < -10 or 3m > 9